OpenStudy (faman39):

Zach borrowed $4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved $4,720 to pay back to his grandfather. At what interest rate did Zach borrow the money? 8% 9% 15% 42%

OpenStudy (ajprincess):

Let the interest rate be x. Additional payment for 2 yrs=4720-4000 =720 So for one year=720/2 =360 x=(Additional payment for one year/Total payment)*100 =(360/4000)*100

OpenStudy (anonymous):